OpenStudy (anonymous):

An airplane is flying horizontally with a velocity of 50.0 m/s and an altitude of 225m when it releases a care package. if we ignore air resistance, how many seconds will it take the care package to fall to the ground.

OpenStudy (anonymous):

Please i really need help with this.

OpenStudy (anonymous):

four and half second

OpenStudy (anonymous):

Thank you but how do you set up the equation needed to solve for time, I have a test coming up soon and I want to get this down.

OpenStudy (anonymous):

Ok..let me tell you

OpenStudy (anonymous):

S=Vt from this equation we can ... t=S/V t=225/50 t=4.5 seconds

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

50 m/s is just for horizontal velocity

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

what is the vertical velocity

OpenStudy (anonymous):

it has no direction it is just 50.0 m/s

OpenStudy (anonymous):

i know but the question said that airplane move horizontally

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

there is no vertical stated but, it says the altitude is 225m

OpenStudy (anonymous):

you should know that this is like a projectile motion

OpenStudy (anonymous):

idk why but it't not in the problem

OpenStudy (anonymous):